OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)?

OpenStudy (anonymous):

find slope using m=(y2-y1)/(x2-x1) then plug it in y-y1=m(x-x1)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

y-11/x-1=11-3/1-(-1) y-4x-7=0